I'm having a problem finding the slope -10x+6y=-30
Answer by Alan3354(69443)

Solve for y
-10x+6y=-30
-5x + 3y = -15
3y = 5x - 15
y = (5/3)x - 5
---------
The slope is the coefficient of x
m = 5/3
That's all.